OpenStudy (anonymous):

f(x) = |3x + 3| in piecewise form?

OpenStudy (anonymous):

\[|x| = \left\{\begin{array}{rcc} -x & \text{if} & x <0 \\ x& \text{if} & x \geq 0 \end{array} \right. \] now replace \(x\) by \(3x+3\) everywhere you see it, and see what you get

OpenStudy (anonymous):

thanks!